if x is an integer chosen at random from the set {2,11,19,31,51} and y is chosen at random from the set {1,7,18,19,31}.What is the probability that the product xy is a prime factor?
Answer by stanbon(75887)

if x is an integer chosen at random from the set {2,11,19,31,51} and y is chosen at random from the set {1,7,18,19,31}.What is the probability that the product xy is a prime product ?
--------
Prime products:: 1*2 , 1*11, !*19, 1*31
None of the other products will be prime.
----
P(xy is prime) = 4/(5*5) = 4/25
--------------------
Cheers,
Stan H.